What is the length of the diagnol of a coordinate plane with x- and y-axes labled 10 to -10.
Answer by algebrahouse.com(1659)

"What is the length of the diagnol of a coordinate plane with x- and y-axes labled 10 to -10."

A right triangle would be created with legs of 20 and 20, and a diagonal of x.

Pythagorean Theorem:
a^2 + b^2 = c^2 {a and b are the legs and c is the hypotenuse (across from the right angle)}

20^2 + 20^2 = c^2 {substituted into the pythagorean theorem}
400 + 400 = c^2 {squared the 20's}
800 = c^2 {combined like terms on left}
c = 28.28 {took square root of both sides}

the diagonal is approximately 28.28
www.algebrahouse.com